A Rectangular Persian Carpet has a perimeter of 164 inches. The length of the carpet is 20 inches more than the width. What are the dimensions of the carpet?
Answer by checkley77(12844)

L=W+20
P=2L+2W
164=2(W+20)+2W
164=2W+40+2W
164=4W+40
4W=164-40
4W=124
W=124/4
W=31 IN. ANS. FOR THE WIDTH.
L=31+20=51 IN. ANS. FOR THE LENGTH.
PROOF:
164=2*51+2*31
164=102+62
164=164